# Efficient sequential Bayesian inference for state-space epidemic models using ensemble data assimilation

### v1

@@ Abstract
Estimating latent epidemic states and model parameters from partially observed, noisy data remains a major challenge in infectious disease modeling. State-space formulations provide a coherent probabilistic framework for such inference, yet fully Bayesian estimation is often computationally prohibitive because evaluating the observed-data likelihood requires integration over all latent trajectories. The Sequential Monte Carlo squared (SMC2) algorithm offers a principled approach for joint state and parameter inference, combining an outer SMC sampler over parameters with an inner particle filter that estimates the likelihood up to the current time point. Despite its theoretical appeal, this nested particle filter imposes substantial computational cost, limiting routine use in near-real-time outbreak response. We propose Ensemble SMC2 (eSMC2), a scalable variant that replaces the inner particle filter with an Ensemble Kalman Filter (EnKF) to approximate the incremental likelihood at each observation time. While this substitution introduces bias via a Gaussian approximation, we mitigate finite-sample effects using an unbiased Gaussian density estimator and adapt the EnKF for epidemic data through state-dependent observation variance. This makes our approach particularly suitable for overdispersed incidence data commonly encountered in infectious disease surveillance. Simulation experiments with known ground truth and an application to 2022 United States (U.S.) monkeypox incidence data demonstrate that eSMC2 achieves substantial computational gains while producing posterior estimates comparable to SMC2. The method accurately recovers latent epidemic trajectories and key epidemiological parameters, providing an efficient framework for sequential Bayesian inference from imperfect surveillance data.
@@ 1 Introduction
Accurate real-time inference of epidemic dynamics is essential for situational awareness, forecasting, and evaluating control interventions during infectious disease outbreaks. Public health decisions often depend on rapidly updating estimates of key quantities such as the time-varying reproduction number, incubation period, and infectious duration. Epidemic surveillance data, however, are often noisy, incomplete, and reported at irregular intervals, making the recovery of latent infection processes a challenging inverse problem: unobserved epidemic states must be inferred from indirect, partial observations of incidence or prevalence (Golightly et al. 2023, Whitaker et al. 2025). This naturally leads to a state-space modeling framework, in which latent epidemic states evolve according to mechanistic transmission dynamics, such as the Susceptible–Exposed–Infectious–Removed (SEIR) model, while observations provide only indirect information about the underlying epidemic process.
@@
State-space models (SSMs) have become central to modern infectious disease modeling because they explicitly account for both stochastic transmission and observational uncertainty (Dureau et al. 2013, Birrell et al. 2018). Within this framework, Bayesian inference offers a principled approach to jointly estimate latent epidemic trajectories and epidemiological parameters while fully quantifying uncertainty. Yet, fully Bayesian inference for SSMs remains computationally demanding: the observed-data likelihood is typically intractable, and standard Monte Carlo techniques may be infeasible for large-scale or real-time applications.
@@
Particle filtering and related methods have made Bayesian state estimation tractable in epidemic modeling, with applications ranging from Influenza, Ebola, COVID-19, and more recently, Monkeypox (Dureau et al. 2013, Funk et al. 2018, Ghosh et al. 2023, Papageorgiou & Kolias 2024). In particular, Particle Markov Chain Monte Carlo (PMCMC) (Andrieu et al. 2010) enables exact Bayesian inference by combining MCMC with unbiased likelihood estimates from a particle filter (PF). However, PMCMC can become prohibitively expensive because it requires running a full particle filter at each MCMC iteration. Moreover high variance of likelihood estimates can lead to poor mixing and increased Monte Carlo error, particularly when the data provide limited information about the parameters. The Sequential Monte Carlo Squared (SMC2) algorithm (Chopin et al. 2013) extends PMCMC by maintaining a population of parameter particles, each associated with an internal particle filter for state inference and likelihood evaluation. Parameter particles are reweighted using the estimated likelihoods, and a resample–move step with a PMCMC mutation kernel mitigates degeneracy and maintains diversity. This nested structure enables sequential Bayesian updating as new data arrive and has been applied successfully in epidemic modeling (Golightly & Kypraios 2018, Temfack & Wyse 2025), but its computational cost grows with particle number, and weight degeneracy remains a concern (Snyder et al. 2008). Although Golightly & Kypraios (2018) demonstrated that using the Auxiliary Particle Filter (APF, Pitt & Shephard 1999) within SMC2 can improve particle selection and reduce the number of particles required compared to a standard bootstrap particle filter (BPF, Gordon et al. 1993), it comes with some trade-offs. The APF requires additional computation and involves more elaborate weighting and proposal steps.
@@
As a computationally efficient and practically attractive alternative, the Ensemble Kalman Filter (EnKF) (Evensen 1994) has demonstrated strong performance in data assimilation for near-Gaussian systems. Compared with PF, the EnKF can achieve reliable results with substantially fewer particles (also referred to as ensemble members), since it updates the ensemble through linear transformations rather than importance resampling, thereby mitigating weight degeneracy. In practice, this often allows the EnKF to outperform particle filters in moderately nonlinear settings where the latter suffer from sample depletion (Morzfeld et al. 2017). The method has also gained traction in epidemiological inference, where it is used to track infection dynamics and estimate time-varying transmission parameters (Mitchell & Arnold 2021, Lal et al. 2021, Sun et al. 2023). Its efficiency and scalability make it particularly appealing for sequential updating as new surveillance data become available. However, most existing epidemic applications employ an augmented EnKF to jointly estimate states and parameters (Lal et al. 2021, Abbas et al. 2025). While effective for combined inference, this approach can be limited in flexibility, potentially struggling to adapt to changes in transmission dynamics. Since parameters are typically treated as additional static states, their artificial evolution (e.g., via random-walk perturbations) can lead to filter degeneracy or drift over time, a well-known issue shared by most Bayesian augmented filtering methods (Stroud et al. 2018).
@@
Several hybrid approaches have been developed to combine the scalability of the EnKF with the inferential rigor of full Bayesian methods. For instance, Stroud et al. (2018) proposed a two-layer filter pairing an EnKF for states with a particle filter for parameters, while Katzfuss et al. (2020) suggested replacing the particle filter entirely with an EnKF in the particle MCMC (pMCMC) framework, an approach that significantly reduces computational cost at the expense of a small approximation bias. Extending this idea, Drovandi et al. (2022) introduced the Ensemble MCMC (eMCMC) algorithm, which systematically corrects the bias in the EnKF likelihood using an unbiased Gaussian density estimator and an early-rejection scheme to improve efficiency. Their results showed that EnKF-based likelihoods can achieve accurate posterior inference in moderately nonlinear systems with orders-of-magnitude speedups compared to particle filters.
@@
In this work, we propose Ensemble SMC2 (eSMC2), a novel algorithm that integrates the EnKF into the inner layer of the SMC2 framework to enable fully sequential Bayesian inference over both latent states and parameters. By replacing the particle filter with an EnKF-based likelihood, eSMC2 substantially reduces computational cost while maintaining posterior accuracy. To mitigate the bias inherent in standard EnKF likelihood approximations, we incorporate the unbiased Gaussian likelihood correction employed by Drovandi et al. (2022), which reduces sensitivity to the number of ensemble members. While we cannot provide formal bias quantification, prior studies indicate that EnKF-based likelihoods yield stable and accurate inference in moderately nonlinear systems (Khalil et al. 2015, Katzfuss et al. 2020, Drovandi et al. 2022). We further employ a variant inspired by the Poisson Kalman Filter (Ebeigbe et al. 2020), making the approach suitable for observations typical in disease surveillance and relaxing the Gaussian noise assumptions of standard EnKF methods. We demonstrate its performance on a diffusion-driven SEIR model with incidence data aggregated over fixed reporting intervals, showing that eSMC2 delivers accurate and scalable inference for time-varying transmission dynamics and key epidemiological parameters using both synthetic and real datasets.
@@
During the final stages of preparing this manuscript, we became aware of the independent preprint by Golightly et al. (2025), which also investigates replacing the particle filter in SMC2 with an EnKF. Their Nested EnKF focuses primarily on general state-space models, high-dimensional settings, and delayed-acceptance acceleration strategies. In contrast, our contribution is motivated by epidemic inference, introduces an unbiased Gaussian likelihood correction, and develops a state-dependent variance adaptation specifically tailored to count-based surveillance data. These complementary developments address different aspects of the broader challenge of scalable Bayesian inference for state-space models, with our work specifically advancing practical tools for near-real-time epidemic monitoring. We concur with Golightly et al. (2025) on the strong performance of this approach.
@@ 2.1 SEIR model in a Bayesian state-space framework
We model the spread of infectious diseases using a stochastic SEIR framework that explicitly tracks both latent epidemic states and the time-varying transmission rate. Let the latent state at time $t$ be $x_{t}=(S_{t},E_{t},I_{t},R_{t},Z_{t},\log(\beta_{t}))^{\top}$, where $S_{t},E_{t},I_{t},R_{t}$ denote the numbers of susceptible, exposed (infected but not yet infectious), infectious, and and removed (recovered or deceased) individuals, respectively. The variable $Z_{t}$ is not a dynamical compartment but an auxiliary quantity representing the cumulative incidence over each reporting interval, included for convenience to directly match the observation model. The variable $\log(\beta_{t})$ is incorporated into the latent state to allow inference on the time-varying transmission rate. The latent dynamics follow the system
@@
where $N=S_{t}+E_{t}+I_{t}+R_{t}$ is the constant population size, and $1/\alpha$ and $1/\gamma$ denote the average incubation and infectious periods, respectively. The transmission rate $\beta_{t}$ evolves according to a Brownian motion on the log scale with volatility parameter $\nu_{\beta}$, driven by a standard Wiener process $B_{t}$. This stochastic formulation allows for smooth, time-varying changes in transmissibility, for example due to interventions or behavioral shifts (Dureau et al. 2013, Cazelles et al. 2018). The effective reproduction number is defined as $R_{\text{eff}}(t)=\beta_{t}S_{t}/(\gamma N)$. To simulate sample trajectories, the continuous-time system (1) is propagated forward using a discrete-time approximation. Let $\theta$ denote the vector of unknown model parameters governing both the latent dynamics in (1) and the observation model described below. The latent state is updated according to
@@
The forward process $p(.|x_{t-1},\theta)$ updates $S_{t},E_{t},I_{t},R_{t},$ and $Z_{t}$ deterministically via a forward Euler scheme, while $\log(\beta_{t})$ evolves stochastically using the Euler–Maruyama method (Kloeden 2011). Formally, the deterministic updates are represented as a Dirac delta distribution within the transition density $p(x_{t}|x_{t-1},\theta)$, whereas the stochasticity of $\log(\beta_{t})$ introduces process noise. Details on the discretization and transition density are provided in Section A.1 of Appendix A.
@@
$y_{t}|x_{t},\theta\sim\text{Poisson}(\rho Z_{t})\quad\text{or}\quad y_{t}|x_{t},\theta\sim\text{NegBin}\big(\mu=\rho Z_{t},\penalty 10000\ \sigma^{2}=\rho Z_{t}+\phi(\rho Z_{t})^{2}\big),$ (3)
where $\rho\in(0,1]$ is the reporting fraction, informed by expert knowledge. Specifying $\rho$ helps mitigate practical non-identifiability issues highlighted in Cazelles et al. (2018), since decreases in observed cases could otherwise be attributed equally to a lower reporting rate, a reduced transmission rate, or a combination of both. Practical considerations regarding the structural identifiability of parameters in such models were also discussed in Temfack & Wyse (2025). In the Negative Binomial model, larger values of $\phi$ correspond to stronger overdispersion, while the Poisson model is recovered in the limit $\phi\to 0$.
Together, Eqs. (2)–(3) define our complete state-space model (Birrell et al. 2018). Let $x_{0:t}=\{x_{0},\ldots,x_{t}\}$ denote the latent states up to time $t$. Prior knowledge about $\theta$ and initial state $x_{0}$ is encoded in the distributions $p(\theta)$ and $p(x_{0}|\theta)$, respectively. The joint posterior of states and parameters given the observations is then
@@
For nonlinear state-space models, Eqs. (5) and (2.1) cannot be solved analytically and are generally intractable. Consequently, practical inference relies on numerical approximations such as particle filtering or EnKF-based methods to generate samples from the filtering distribution or to estimate the marginal likelihood. In this work, motivated by the computational efficiency and scalability considerations discussed in the introduction, we focus on EnKF-based inference.
@@ 2.2 Ensemble Kalman Filter
We now describe the ensemble Kalman filter (EnKF) (Evensen 1994), a sequential inference method for state-space models with nonlinear latent dynamics. While the EnKF was originally developed for systems with Gaussian observations, many real-world datasets, including those arising in epidemic surveillance often deviate from this assumption and may exhibit overdispersion. To accommodate such data, we follow the approach of Ebeigbe et al. (2020) and extend it by constructing observation noise terms consistent with the observation models in Eq. (3).
The EnKF alternates between prediction and update operations at each time step to sequentially refine the latent state estimate based on incoming observations (Figure 1). During the prediction step, the ensemble is propagated forward through the system dynamics, while in the update step, the ensemble is adjusted using the most recent observation.
@@
$p(x_{t}|y_{1:t-1},\theta)\approx\mathcal{N}\big(x_{t};\penalty 10000\ \widehat{\mu}_{t|t-1},\penalty 10000\ \widehat{\Sigma}_{t|t-1}\big),$ (9)
@@
A key adaptation for epidemic data is the specification of the observation error variance $V_{t}$. Classical Kalman filtering assumes Gaussian observations with constant variance (or at least does not depend on the latent state), an assumption often violated for epidemiological models, where the variability of reported incidence typically scales with its magnitude (Ebeigbe et al. 2020). A principled approach is to set $V_{t}$ to match the expected conditional variance of the observation given the latent state $\mathrm{E}\big[\mathrm{Var}[y_{t}|x_{t}]\big]$, as implied by the observation model (3). In practice we approximate the expectation above using the forecast ensemble. We define the ensemble estimate of the observation error variance by
@@
$V_{t}=\max\big(\eta,V_{t}^{N_{x}}\big),$ (11)
where in our implementation we set $\eta=0.1$, small relative to the typical magnitude of $V_{t}^{N_{x}}$. We found that results are insensitive to this choice provided $\eta<1$. Section A.2 of Appendix A provides a detailed justification for this specification. This specification allows the EnKF update to capture key features of the observation distribution (3), though it can overestimate uncertainty when case counts are extremely low. Such low-incidence scenarios are not the primary focus of our study, and alternative sequential methods, such as the Lifebelt particle filter (Corbella et al. 2024), may be preferred in those cases. A similar treatment of adaptive variance specification is discussed in (Section 3.2, Asfaw et al. 2024).
@@
$x_{t}^{(i)}=x_{t}^{(f,i)}+K_{t}\left(y_{t}+v_{t}^{(i)}-Hx_{t}^{(f,i)}\right),\quad v_{t}^{(i)}\sim\mathcal{N}(0,V_{t}),$ (12)
@@
$K_{t}=\widehat{\Sigma}_{t|t-1}H^{\top}\left(H\widehat{\Sigma}_{t|t-1}H^{\top}+V_{t},\right)^{-1}.$ (13)
@@
$p(y_{t}|y_{1:t-1},\theta)\approx\mathcal{N}\Big(y_{t};\penalty 10000\ H\widehat{\mu}_{t|t-1},\penalty 10000\ H\widehat{\Sigma}_{t|t-1}H^{\top}+V_{t}\Big).$ (14)
Only the key components of the EnKF algorithm for state inference are presented above; for a comprehensive overview, see Katzfuss et al. (2020). For linear–Gaussian SSMs, the likelihood approximation in (14) converges to the true likelihood as $N_{x}\to\infty$, but for finite ensembles it remains biased due to sampling variability in the forecast mean and covariance (Drovandi et al. 2022). Importantly, even if the forecast ensemble is exactly Gaussian, the estimated density $\mathcal{N}(y_{t};\penalty 10000\ H\widehat{\mu}_{t|t-1},\penalty 10000\ H\widehat{\Sigma}_{t|t-1}H^{\top}+V_{t})$ is not an unbiased estimate of the true likelihood. This occurs because the empirical normal density $\mathcal{N}(y;\penalty 10000\ \mu_{{}_{N_{x}}},\Sigma_{{}_{N_{x}}})$, computed from the finite-ensemble mean $\mu_{{}_{N_{x}}}$ and covariance $\Sigma_{{}_{N_{x}}}$, is biased estimator of the true density $\mathcal{N}(y;,\mu,\Sigma)$ for any finite $N_{x}$.
Nevertheless, empirical results in the literature (Khalil et al. 2015, Stroud et al. 2018, Drovandi et al. 2022) show that the EnKF likelihood performs remarkably well for inference in both linear and nonlinear systems, with posterior estimates only weakly dependent on the ensemble size (Drovandi et al. 2022). To mitigate the finite-sample bias, we adopt the unbiased Gaussian density estimator of Ghurye & Olkin (1969), later employed by Price et al. (2018) and Drovandi et al. (2022). This estimator yields an unbiased estimate of $\mathcal{N}(y;\mu,\Sigma)$ provided $N_{x}>d_{y}+3$, where $d_{y}$ is the dimension of $y$. It is defined as
$\widehat{\mathcal{N}}_{\text{unb}}(y;\penalty 10000\ \mu_{{}_{N_{x}}},\Sigma_{{}_{N_{x}}})=$ $\ (2\pi)^{-d_{y}/2}\frac{c(d_{y},N_{x}-2)}{c(d_{y},N_{x}-1)(1-1/N_{x})^{d_{y}/2}}|M_{{}_{N_{x}}}|^{-(N_{x}-d_{y}-2)/2}$

### v2

@@ Abstract
Estimating latent epidemic states and model parameters from partially observed, noisy data remains a major challenge in infectious disease modeling. State-space formulations provide a coherent probabilistic framework for such inference, yet fully Bayesian estimation is often computationally prohibitive because evaluating the observed-data likelihood requires integration over a latent trajectory. The Sequential Monte Carlo squared (SMC2) algorithm offers a principled approach for joint state and parameter inference, combining an outer SMC sampler over parameters with an inner particle filter that estimates the likelihood up to the current time point. Despite its theoretical appeal, this nested particle filter imposes substantial computational cost, limiting routine use in near-real-time outbreak response. We propose Ensemble SMC2 (eSMC2), a computationally efficient variant that replaces the inner particle filter with an Ensemble Kalman Filter (EnKF) to approximate the incremental likelihood at each observation time. While this substitution introduces bias via a Gaussian approximation, we mitigate finite-sample effects using an unbiased Gaussian density estimator and adapt the EnKF for epidemic data through state-dependent observation variance. This makes our approach particularly suitable for overdispersed incidence data commonly encountered in infectious disease surveillance. Simulation experiments with known ground truth and an application to 2022 United States (U.S.) monkeypox incidence data demonstrate that eSMC2 achieves substantial computational gains while producing posterior estimates comparable to SMC2. The method accurately recovers latent epidemic trajectories and key epidemiological parameters, providing an efficient framework for sequential Bayesian inference from imperfect surveillance data.
@@ 1 Introduction
Accurate real-time inference of epidemic dynamics is essential for situational awareness, forecasting, and evaluating interventions during outbreaks. Public health decisions rely on rapidly updating estimates of quantities such as the time-varying reproduction number, incubation period, and infectious duration. Epidemic surveillance data are often noisy, incomplete, and reported irregularly, making the recovery of latent infection processes a challenging inverse problem: the unobserved transmission process must be inferred from partial, indirect observations of incidence or prevalence (Golightly et al., 2023; Whitaker et al., 2025). This motivates a state-space modeling framework, in which latent epidemic states evolve according to mechanistic transmission dynamics, such as the Susceptible–Exposed–Infectious–Removed (SEIR) model, while observations provide indirect information about the underlying epidemic process.
@@
State-space models (SSMs) have become central to modern infectious disease modeling because they explicitly account for stochastic transmission and observational uncertainty (Dureau et al., 2013; Birrell et al., 2018). Within this framework, Bayesian inference provides a principled approach to jointly estimate latent epidemic trajectories and epidemiological parameters. However, fully Bayesian inference for SSMs remains computationally demanding: the observed-data likelihood is typically intractable, and standard Monte Carlo methods may be impractical for large-scale or real-time applications.
@@
Particle filtering and related methods have made Bayesian state estimation tractable in epidemic modeling, with applications to Influenza, Ebola, COVID-19, and more recently, Monkeypox (Dureau et al., 2013; Funk et al., 2018; Ghosh et al., 2023; Papageorgiou and Kolias, 2024). Particle Markov Chain Monte Carlo (PMCMC) (Andrieu et al., 2010) enables exact Bayesian inference by combining MCMC with unbiased likelihood estimates from a particle filter (PF), but it can be computationally expensive since a full PF must be run at each MCMC iteration. Likelihood estimates may also have high variance, leading to poor mixing and increased Monte Carlo error, especially when the data provide limited information about parameters. Sequential Monte Carlo Squared (SMC2) (Chopin et al., 2013) extends PMCMC by maintaining a population of parameter particles, each with an internal PF for state inference and likelihood evaluation. Parameter particles are reweighted, and a resample–move step with a PMCMC mutation kernel mitigates degeneracy and maintains diversity. This nested structure allows sequential Bayesian updating as new data arrive and has been applied successfully in epidemic modeling (Golightly and Kypraios, 2018; Temfack and Wyse, 2025), but computational cost grows with particle number and weight degeneracy remains a concern (Snyder et al., 2008). Using the Auxiliary Particle Filter (APF) (Pitt and Shephard, 1999) within SMC2 can improve particle selection and reduce the number of particles required (Golightly and Kypraios, 2018), but it adds computational complexity due to more elaborate weighting and proposal steps. Likelihood-free approaches, notably Approximate Bayesian Computation (ABC), offer an alternative when the likelihood is fully intractable (Li et al., 2025). However, ABC typically requires many model simulations and careful choice of summary statistics and thresholds, which can make real-time application challenging.
@@
As a computationally efficient and appealing method, the Ensemble Kalman Filter (EnKF, Evensen, 1994) has demonstrated strong performance in data assimilation for near-Gaussian systems. Compared with PF, the EnKF can achieve reliable results with substantially fewer particles (also referred to as ensemble members), since it updates the ensemble through linear transformations rather than importance resampling, thereby mitigating weight degeneracy. In practice, this often allows the EnKF to outperform particle filters in moderately nonlinear settings where the latter suffer from sample depletion (Morzfeld et al., 2017). In epidemiological applications, the EnKF is used to track infection dynamics and estimate time-varying transmission parameters (Mitchell and Arnold, 2021; Lal et al., 2021; Sun et al., 2023). Most implementations employ an augmented EnKF to jointly estimate states and parameters (Lal et al., 2021; Abbas et al., 2025). While effective, this approach can be less flexible in adapting to changing transmission dynamics, since parameters are treated as additional static states and their artificial evolution (e.g., via random-walk perturbations) may lead to filter degeneracy or drift over time, a well-known issue shared by most Bayesian augmented filtering methods (Stroud et al., 2018).
@@
Several hybrid approaches have been developed to combine the scalability of the EnKF with the inferential rigor of full Bayesian methods. For instance, Stroud et al. (2018) proposed a two-layer filter pairing an EnKF for states with a particle filter for parameters, while Katzfuss et al. (2020) suggested replacing the particle filter with an EnKF within the particle MCMC (PMCMC) framework, an approach that substantially reduces computational cost at the expense of a small approximation bias. Extending this idea, Drovandi et al. (2022) introduced the Ensemble MCMC (eMCMC) algorithm, which corrects the bias in the EnKF likelihood using an unbiased Gaussian density estimator together with an early-rejection scheme to improve efficiency. Their results showed that EnKF-based likelihoods can achieve accurate posterior inference in moderately nonlinear systems with orders-of-magnitude speedups compared to particle filters.
@@
In this work, we propose Ensemble SMC2 In this work, we propose Ensemble SMC2 (eSMC2), a novel algorithm that integrates the EnKF into the inner layer of the SMC2 framework to enable fully sequential Bayesian inference over both latent states and parameters. This approach substantially reduces computational cost while maintaining accurate posterior inference. To mitigate the bias inherent in the standard EnKF likelihood estimate, we incorporate the unbiased Gaussian likelihood correction employed by Drovandi et al. (2022), which reduces sensitivity to the number of ensemble members. While we cannot provide formal bias quantification, prior studies indicate that EnKF-based likelihoods yield stable and accurate inference in moderately nonlinear systems (Khalil et al., 2015; Katzfuss et al., 2020; Drovandi et al., 2022). We further employ a variant inspired by the Poisson Kalman Filter (Ebeigbe et al., 2020), making the approach suitable for observations typical in disease surveillance and relaxing the Gaussian noise assumptions of standard EnKF methods. We demonstrate its performance on a diffusion-driven SEIR model with incidence data aggregated over fixed reporting intervals, showing that eSMC2 delivers an accurate estimate of the time-varying transmission dynamics and key epidemiological parameters using both synthetic and real datasets.
@@
During the final stages of this work, we became aware of the independent preprint by Golightly et al. (2025), which also replaces the particle filter in SMC2 with an EnKF. Their Nested EnKF focuses on general state-space models and delayed-acceptance strategies, whereas our approach is motivated by epidemic inference and incorporates an unbiased Gaussian likelihood correction with variance adaptation for count-based surveillance data. These complementary developments address different aspects of Bayesian inference for state-space models, and we concur with (Golightly et al., 2025) on the strong performance of this approach.
@@ 2.1 SEIR model in a Bayesian state-space framework
We model the spread of infectious diseases using a stochastic SEIR framework that explicitly tracks both latent epidemic states and the time-varying transmission rate. While we present the SEIR model here as a concrete example, the methodology is applicable to a wide range of compartmental epidemic models. Let the latent state at time $t$ be $x_{t}=(S_{t},E_{t},I_{t},R_{t},Z_{t},\log(\beta_{t}))^{\top}$, where $S_{t},E_{t},I_{t},R_{t}$ denote the numbers of susceptible, exposed (infected but not yet infectious), infectious, and and removed (recovered or deceased) individuals, respectively. The variable $Z_{t}$ is not a dynamical compartment but an auxiliary quantity representing the cumulative incidence over each reporting interval, included for convenience to directly match the observation model. The variable $\log(\beta_{t})$ is incorporated into the latent state to allow inference on the time-varying transmission rate. The latent dynamics follow the system
@@
where $N=S_{t}+E_{t}+I_{t}+R_{t}$ is the constant population size, and $1/\alpha$ and $1/\gamma$ denote the average incubation and infectious periods, respectively. The transmission rate $\beta_{t}$ is modeled as a stochastic process on the log scale, following a Brownian motion with volatility $\nu_{\beta}$ and driven by a standard Wiener process $B_{t}$. This allows $\beta_{t}$ to vary over time in response to behavioral changes, public health interventions, or seasonal effects (Dureau et al., 2013; Cazelles et al., 2018). The parameter $\nu_{\beta}$ controls the magnitude of these fluctuations: larger values permit faster changes, while smaller values constrain variation. This formulation enables the model to capture unobserved temporal variability in transmission intensity throughout an epidemic. The effective reproduction number is defined as $R_{\text{eff}}(t)=\beta_{t}S_{t}/(\gamma N)$. To simulate sample trajectories, the continuous-time system (1) is propagated forward using a discrete-time approximation. Let $\theta$ denote the vector of unknown model parameters governing both the latent dynamics in (1) and the observation model described below. The latent state is updated according to
@@
The forward process $p(.|x_{t-1},\theta)$ updates $S_{t},E_{t},I_{t},R_{t},$ and $Z_{t}$ deterministically via a forward Euler scheme, while $\log(\beta_{t})$ evolves stochastically using the Euler–Maruyama method (Kloeden, 2011). Formally, the deterministic updates are represented as a Dirac delta distribution within the transition density $p(x_{t}|x_{t-1},\theta)$, whereas the stochasticity of $\log(\beta_{t})$ introduces process noise. Details on the discretization and transition density are provided in Appendix A.
@@
$y_{t}|x_{t},\theta\sim\text{Poisson}(\rho Z_{t})\quad\text{or}$ (3a)
$y_{t}|x_{t},\theta\sim\text{NegBin}\big(\mathrm{mean}=\rho Z_{t},~\mathrm{var}=\rho Z_{t}+\phi(\rho Z_{t})^{2}\big),$ (3b)
The Negative Binomial is parameterized by its mean $\rho Z_{t}$ and variance $V_{t}=\mathrm{Var}[y_{t}|x_{t}]=\rho Z_{t}+\phi(\rho Z_{t})^{2}$, where $\phi>0$ is the overdispersion parameter; the Poisson model, with $V_{t}=\mathrm{Var}[y_{t}|x_{t}]=\rho Z_{t}$, is recovered as $\phi\to 0$. This mean-variance parameterization is used directly in the ensemble-based observation variance described in Section 2.2 below, ensuring consistency between the observation model and the EnKF update step. $\rho\in(0,1]$ is the reporting fraction, informed by expert knowledge. Specifying $\rho$ helps mitigate practical non-identifiability issues highlighted in Cazelles et al. (2018), since decreases in observed cases could otherwise be attributed equally to a lower reporting rate, a reduced transmission rate, or a combination of both. Practical considerations regarding the structural identifiability of parameters in such models were also discussed in Temfack and Wyse (2025).
Together, Eqs (2)–(3) define our complete state-space model (Birrell et al., 2018). Let $x_{0:t}=\{x_{0},\ldots,x_{t}\}$ denote the latent states up to time $t$. Prior knowledge about $\theta$ and initial state $x_{0}$ is encoded in the distributions $p(\theta)$ and $p(x_{0}|\theta)$, respectively. The joint posterior of states and parameters given the observations is then
@@
For nonlinear state-space models, Eqs (5) and (2.1) cannot be solved analytically and are generally intractable. Consequently, practical inference relies on numerical approximations such as particle filtering or EnKF-based methods to generate samples from the filtering distribution or to estimate the marginal likelihood. In this work, motivated by the computational efficiency and scalability considerations discussed in the introduction, we focus on EnKF-based inference.
@@ 2.2 Ensemble Kalman Filter
We now describe the ensemble Kalman filter (EnKF, Evensen, 1994), a sequential inference method for state-space models with nonlinear latent dynamics. While the EnKF was originally developed for systems with Gaussian observations, many real-world datasets, including those arising in epidemic surveillance often deviate from this assumption and may exhibit overdispersion. To accommodate such data, we follow the approach of Ebeigbe et al. (2020) and extend it by constructing observation noise terms consistent with the observation models in Eq (3).
The EnKF alternates between prediction and update operations at each time step to sequentially refine the latent state estimate based on incoming observations (Fig 1). During the prediction step, the ensemble is propagated forward through the system dynamics, while in the update step, the ensemble is adjusted using the most recent observation.
[…]
@@
$p(x_{t}|y_{1:t-1},\theta)\approx\mathcal{N}\big(x_{t};~\widehat{\mu}_{t|t-1},~\widehat{\Sigma}_{t|t-1}\big),$ (9)
@@
A key adaptation for epidemic data is the specification of the observation error variance $V_{t}$. Classical Kalman filtering assumes Gaussian observations with constant variance (or at least does not depend on the latent state), an assumption often violated for epidemiological models, where the variability of reported incidence typically scales with its magnitude (Ebeigbe et al., 2020). A principled approach is to set $V_{t}$ to match the expected conditional variance of the observation given the latent state $\mathrm{E}\big[\mathrm{Var}[y_{t}|x_{t}]\big]$, as implied by the observation model (3). In practice, we approximate the expectation above using the forecast ensemble. We define the ensemble estimate of the observation error variance by
@@
$\widehat{V}_{t}=\max\big(\eta,V_{t}^{N_{x}}\big),$ (11)
where in our implementation we set $\eta=0.1$, small relative to the typical magnitude of $V_{t}^{N_{x}}$. We found that results are insensitive to this choice provided $\eta<1$. Appendix A provides a detailed justification for this specification. This specification allows the EnKF update to capture key features of the observation distribution (3), though it can overestimate uncertainty when case counts are extremely low. Such low-incidence scenarios are not the primary focus of our study, and alternative sequential methods, such as the Lifebelt particle filter (Corbella et al., 2024), may be preferred in those cases. A similar treatment of adaptive variance specification is discussed in (Section 3.2, Asfaw et al. (2024)).
@@
$x_{t}^{(i)}=x_{t}^{(f,i)}+K_{t}\left(y_{t}+v_{t}^{(i)}-Hx_{t}^{(f,i)}\right),\quad v_{t}^{(i)}\sim\mathcal{N}(0,\widehat{V}_{t}),$ (12)
@@
$K_{t}=\widehat{\Sigma}_{t|t-1}H^{\top}\left(H\widehat{\Sigma}_{t|t-1}H^{\top}+\widehat{V}_{t},\right)^{-1}.$ (13)
@@
$p(y_{t}|y_{1:t-1},\theta)\approx\mathcal{N}\Big(y_{t};~H\widehat{\mu}_{t|t-1},~H\widehat{\Sigma}_{t|t-1}H^{\top}+\widehat{V}_{t}\Big).$ (14)
Only the key components of the EnKF algorithm for state inference are presented above; for a comprehensive overview, see Katzfuss et al. (2020). For linear–Gaussian SSMs, the likelihood approximation in (14) converges to the true likelihood as $N_{x}\to\infty$, but for finite ensembles it remains biased due to sampling variability in the forecast mean and covariance (Drovandi et al., 2022). Importantly, even if the forecast ensemble is exactly Gaussian, the estimated density $\mathcal{N}(y_{t};~H\widehat{\mu}_{t|t-1},~H\widehat{\Sigma}_{t|t-1}H^{\top}+\widehat{V}_{t})$ is not an unbiased estimate of the true likelihood. This occurs because the empirical normal density $\mathcal{N}(y;~\mu_{{}_{N_{x}}},\Sigma_{{}_{N_{x}}})$, computed from the finite-ensemble mean $\mu_{{}_{N_{x}}}$ and covariance $\Sigma_{{}_{N_{x}}}$, is biased estimator of the true density $\mathcal{N}(y;,\mu,\Sigma)$ for any finite $N_{x}$.
Nevertheless, empirical results in the literature (Khalil et al., 2015; Stroud et al., 2018; Drovandi et al., 2022) show that the EnKF likelihood performs remarkably well for inference in both linear and nonlinear systems, with posterior estimates only weakly dependent on the ensemble size (Drovandi et al., 2022). To mitigate the finite-sample bias, we adopt the unbiased Gaussian density estimator of Ghurye and Olkin (1969), later employed by Price et al. (2018) and Drovandi et al. (2022). This estimator yields an unbiased estimate of $\mathcal{N}(y;\mu,\Sigma)$ provided $N_{x}>d_{y}+3$, where $d_{y}$ is the dimension of $y$. It is defined as
$\widehat{\mathcal{N}}_{\text{unb}}(y;~\mu_{{}_{N_{x}}},\Sigma_{{}_{N_{x}}})=$ $\ (2\pi)^{-d_{y}/2}\frac{c(d_{y},N_{x}-2)}{c(d_{y},N_{x}-1)(1-1/N_{x})^{d_{y}/2}}|M_{{}_{N_{x}}}|^{-(N_{x}-d_{y}-2)/2}$